OpenStudy (anonymous):

I need help in Series Question. The series is : (1/2) + (2/3)x + (3/4)^2x^2 + (4/5)^3x^3 this is test for convergence

OpenStudy (amistre64):

you prolly need to define a rule for the sequence of coefficients

OpenStudy (amistre64):

\[\sum_{n=0}^{\infty}~\frac{n+1}{n+2}x^n\] seems apropriate to play with

OpenStudy (amistre64):

might be reading some of it wring tho

OpenStudy (amistre64):

\[\frac12+\sum_{n=1}^{\infty}~\frac{(n+1)^n}{(n+2)^n}x^n\] might be better

OpenStudy (amistre64):

\[\lim\frac{(n+1+1)^{n+1}x^{n+1}}{(n+1+2)^{n+1}}\frac{(n+2)^n}{(n+1)^nx^{n}}\] \[|x|\lim\frac{(n+2)^{2n+1}}{(n+3)^{n+1}(n+1)^n}\] might be able to determine of the rational expression is a higher degree top to bottom

OpenStudy (amistre64):

as n gets large this simplifies to:\[\frac{n^{2n+1}}{n^{n+1}*n^n}\to1\] if i see it right

OpenStudy (anonymous):

yes and thats where the problem lies... it goes to 1.. when i apply root test x>1 its divergent. x<1 its convergent but what should I do when x=1??

OpenStudy (amistre64):

the coeffs go to 1, not the limit of the function

OpenStudy (amistre64):

|x| < 1, when -1 < x < 1 this is the interval of convergence

OpenStudy (amistre64):

when x >= 1, you either have to try a different appraoch or just be sated that its divergent ....

OpenStudy (anonymous):

ok thanx

OpenStudy (amistre64):

good luck

OpenStudy (anonymous):

i just joined how to give u medals or whatever for your effort in helping??

OpenStudy (amistre64):

there should be a "best respose" button to the right of all my posts. if not, hit refresh. but a simple thank you is worth more to me than some fictitious award :)

OpenStudy (anonymous):

done both :D

OpenStudy (amistre64):

again, good luck ;)